The ratio of a rectangle’s length to the length of a square’s side is 3:1. If the area of the square is 36, and the rectangle’s width is 2, what is the area of the rectangle?
(A) 24 (B) 36 (C) 72 (D) 108 (E) 216

Answer by richwmiller(17219)

rl/sl=3:1=3/1
area of square=s^2=36
sl=6
rl=3*sl=18
area of rectangle=L*W=18*2=36
(B) 36